OpenStudy (migitmack):

A company builds computers. It costs $6,700 to build 10 computers and $12,200 to build 20 computers. Which equation models the cost, C(x), as a linear function of the number of computers built, x? C(x) = 550x −1200 C(x) = 550x + 1200 C(x) = 1200x − 550 C(x) = 1200x + 550

OpenStudy (migitmack):

Any help?

OpenStudy (anonymous):

Given the information, you have two points on your line: (10, 6700) and (20, 12200) where the x-coordinate is computers and the y-coordinate is money! Now lets find a line that crosses these two points! First, lets find the slope: (12200-6700)/(20-10) = 550 Then, to find the y-intercept, you plug one of those points into y=mx+b with the slope that you just found. So: y=mx+b 6700 = 550*10 +b b = 6700 - 5500 = 1200 Now you have your slope and y-intercept, the two things you need to find your line! C(x) = 550x + 1200 =)